# CT-AGRG: Automated Abnormality-Guided Report Generation from 3D Chest CT Volumes

### v7

@@ 6.2 Ablation study
As shown in Table 1, we conduct an ablation study to evaluate each component of our CT-AGRG model. Introducing multi-task classification significantly improves F1-Score from 23.80 to 46.11 ($\Delta$+93.76increase). Augmenting the latent space further enhances performance, increasing Recall from 56.56 to 58.89 ($\Delta$+4.12increase) and F1-Score from 46.11 to 47.22 ($\Delta$+2.41increase). NLG metrics show notable improvements, with BLEU-4 [80] and METEOR [81] scores increasing by 0.089 and 0.069, respectively, indicating enhanced sentence variability while maintaining semantic fidelity to ground truth reports.
@@ 7 Conclusion
In this paper, we have presented an effective method for guided report generation. Our approach, based on conditional description generation for each abnormality, outperforms state-of-the-art methods while offering a high level of control over the generated report. Future work could focus on adapting the approach to require less supervision, leveraging other modalities or region-specific data to further enhance the report generation process.
@@ 8 Acknowledgments
We acknowledge Hamamci et al. 2024 [14] for making the CT-RATE dataset available.

### v8

@@ 6.2 Ablation study
As shown in Table 1, we conduct an ablation study to evaluate each component of our CT-AGRG model. Introducing multi-task classification significantly improves F1-Score from 23.80 to 46.11 ($\Delta$+93.76% increase). Augmenting the latent space further enhances performance, increasing Recall from 56.56 to 58.89 ($\Delta$+4.12% increase) and F1-Score from 46.11 to 47.22 ($\Delta$+2.41% increase). NLG metrics show notable improvements, with BLEU-4 [papineni_bleu_2002] and METEOR [lavie_meteor_2009] scores increasing by 0.089 and 0.069, respectively, indicating enhanced sentence variability while maintaining semantic fidelity to ground truth reports.
@@ 7 Conclusion
In this paper, we have presented an effective method for guided report generation. Our approach, based on conditional description generation for each abnormality, outperforms state-of-the-art methods while offering a high level of control over the generated report. Future work could focus on adapting the approach to require less supervision, leveraging other modalities or region-specific data to further enhance the report generation process. Access CT-AGRG implementation at https://github.com/theodpzz/ct-agrg.
@@ 8 Acknowledgments
We acknowledge Hamamci et al. 2024 [hamamci_foundation_2024] for making the CT-RATE dataset available.